# Unsupervised learning in echo state networks for input reconstruction

### v3

@@ Abstract
Conventional echo state networks (ESNs) require supervised learning to train the readout layer, using the desired outputs as training data. In this study, we focus on input reconstruction (IR), which refers to training the readout layer to reproduce the input time series in its output. We reformulate the learning algorithm of the ESN readout layer to perform IR using unsupervised learning (UL). By conducting theoretical analysis and numerical experiments, we demonstrate that IR in ESNs can be effectively implemented under realistic conditions without explicitly using the desired outputs as training data; in this way, UL is enabled. Furthermore, we demonstrate that applications relying on IR, such as dynamical system replication and noise filtering, can be reformulated within the UL framework. Our findings establish a theoretically sound and universally applicable IR formulation, along with its related tasks in ESNs. This work paves the way for novel predictions and highlights unresolved theoretical challenges in ESNs, particularly in the context of time-series processing methods and computational models of the brain.
keywords:
Echo state networks; noise filtering; reservoir computing; blind source separation; computational neuroscience; unsupervised learning.
[…]
@@ 1 Introduction
Reservoir computing (RC) is an efficient computational framework for time-series processing. Although RC implementations are diverse, models, such as the echo state network (ESN) (Jaeger, 2001b) and the liquid state machine (Maass et al., 2002), share the same structure: fixed input and reservoir layers and a learnable readout layer (Figure 1(a)). Readers may refer to (Tanaka et al., 2019) for a comprehensive review of RC. The learnable readout layer in RC is trained to match the desired output time series. To define such a time series for RC, we need to define a time-series task. In some cases, the input time series itself can also be used as the desired output. In this paper, we refer to this task as input reconstruction.
@@
Input reconstruction (IR) is the simplest yet fundamental task. The RC performance is based on two time-series processing characteristics: preservability and transformability of the input time series within RC (Maass et al., 2002; Jaeger, 2001a). Considering this, IR requires the preservation of the input without transformation; consequently, it is the most basic time-series processing task. Despite its simplicity, IR in RC is sufficient for the learning of a dynamical system that generates an input time series (Lu et al., 2017, 2018; Röhm et al., 2021; Grigoryeva et al., 2023). Here, we focus on IR to investigate the fundamental properties of RC.
@@
IR has been studied not only within the context of RC but also through two major unsupervised learning (UL) approaches: noise filtering (Candy, 2005) and blind source separation (Ghahramani, 2004). Noise filtering methods, such as the Kalman filter (Kalman, 1960) and its nonlinear extensions (see (Daum, 2005) for a review), rely on the prior knowledge of a system model governing the input time series. In noise filtering, this prior knowledge is used to reconstruct the original input time series or system state using noisy observations (Figure 1(b)). In contrast, blind source separation relies solely on statistical properties without assuming a specific system model: this approach includes principal component analysis (Pearson, 1901), independent component analysis (Jutten and Herault, 1991), and slow feature analysis (Wiskott and Sejnowski, 2002). Compared with noise filtering, blind source separation typically provides an approximate form of IR by identifying the permuted and scaled versions of the input signals. This method relies on minimal assumptions, such as invertibility, to reconstruct unobserved source signals obtained from mixed observations (Figure 1(c)).
@@
However, the current standard approach in RC for IR is supervised learning (SL) of the readout layer (Figure 1(a)). Studies on combining RC with noise filtering (e.g., (Han and Wang, 2009; Tsai et al., 2010; Tomizawa and Sawada, 2021; Goswami et al., 2021)) or blind source separation (e.g., (Antonelo and Schrauwen, 2009; Dinh and Hinaut, 2020; Steiner et al., 2023)) have been conducted; however, in those, the unobserved original input time series was used as supervised data for training the readout layer. Although the listed studies are not exhaustive, this reliance on supervised data appears to be common. We believe that this problem formulation is unnatural compared with noise filtering and blind source separation for IR.
@@
The objective of this study is to refine and advance the IR problem formulation in RC. We propose an IR formulation based on UL. Specifically, we propose a method to replace the availability assumption of an unobserved original input time series with prior knowledge of RC parameters that satisfy certain conditions (Figure 1(d)). The proposed method enables us to formulate IR independently of a specific input time series, allowing for a canonical representation of IR in RC. To illustrate the utility of this novel formulation, we develop algorithms for dynamical system replication and noise filtering that do not explicitly rely on the unobserved original input time series. In addition, we conduct numerical experiments to demonstrate the feasibility of the proposed formulation.
@@
This paper is organized as follows. In Section 2, we define IR in ESNs and propose its reformulation based on UL. In Section 3, we demonstrate that dynamical system replication and noise filtering in ESNs can be theoretically achieved via IR, thus reformulated as UL tasks. Additionally, we provide numerical examples illustrating both tasks solved using the proposed method. In Section 4, we discuss the implications of our findings, highlighting how IR, dynamical system replication, and noise filtering in ESNs are fundamentally redefined under the UL framework. Finally, we present our conclusions in Section 5.
@@ 2 Unsupervised input reconstruction
In this section, we formalize unsupervised learning (UL) for input reconstruction (IR) in RC. Initially, we introduce ESN (Jaeger, 2001b) as a numerically implementable and mathematically tractable RC model. Next, we formalize IR in ESNs. Finally, we describe the mathematical conditions that enable UL for IR in ESNs.
@@ 2.1 Echo state network
$r_{t+1}=g(r_{t},d_{t}):=\sigma\left[Ar_{t}+Bd_{t}\right],$ (1)
where $\sigma:\mathbb{R}^{n_{\text{r}}}\to\mathbb{R}^{n_{\text{r}}}$, $A\in\mathbb{R}^{n_{\text{r}}\times n_{\text{r}}}$, and $B\in\mathbb{R}^{n_{\text{r}}\times n_{\text{in}}}$. Function $\sigma$ and matrices $A$ and $B$ are fixed for each ESN. Function $\sigma$ is referred to as the activation function, and matrices $A$ and $B$ are referred to as the reservoir layer and the input layer, respectively.
@@
The standard approach to learning $W$ is the least squares method which requires the desired output data $o_{t}$. The least squares solution for the readout layer parameter that minimizes $\|Wr_{t}-o_{t}\|^{2}$ is
$W_{\mathcal{O}}:=\mathcal{O}_{1,T}\mathcal{R}_{1,T}^{+}\in\mathbb{R}^{n_{\text{out}}\times n_{\text{r}}},$ (2)
@@ 2.2 Input reconstruction
In IR, we consider the case where the desired ESN output is equal to its input, i.e., $o_{t}=d_{t}(n_{\text{out}}=n_{\text{in}})$ for the entire time $t$. Then, using the standard approach introduced in (2), the solution for IR is
$W_{\mathcal{D}}:=\mathcal{D}_{1,T}\mathcal{R}_{1,T}^{+}\in\mathbb{R}^{n_{\text{out}}\times n_{\text{r}}}.$ (3)
Notation $W_{\mathcal{D}}$ emphasizes that its calculation requires samples obtained from the original input time series $\mathcal{D}_{1,T}$. Thus, the standard approach for IR in ESN assumes that samples obtained from the original input time series $\mathcal{D}_{1,T}$ are available during the training of the readout layer. Nevertheless, its objective is to reconstruct the original input during testing. We address the distinction in the problem settings between training and testing by introducing a training method for the readout layer that does not explicitly rely on the original input.
@@ 2.3 Unsupervised learning for input reconstruction
2.3.1 Main statements
The most vital key point of our method is that the original input $d_{t}$ can be expressed using reservoir state $r_{t}$ under the invertibility assumptions on $\sigma$ and $B$. Specifically, we have the following lemma:
@@ 2.3.1 Main statements
$d_{t}=B^{+}\left[\sigma^{-1}(r_{t+1})-Ar_{t}\right]$ for all $t\in\mathbb{Z}$ if $\sigma$ is invertible and $\text{rank}(B)=n_{\text{in}}$.
@@
Solve (1) for $d_{t}$ using $B^{+}B=I$. ∎
@@
$\mathcal{D}_{1,T}=B^{+}\left[\sigma^{-1}(\mathcal{R}_{2,T+1})-A\mathcal{R}_{1,T}\right]$ for all $t\in\mathbb{Z}$ if $\sigma$ is invertible and $\text{rank}(B)=n_{\text{in}}$.
@@
The least squares solution for IR in ESNs is expressed as
$W_{\mathcal{R}}=B^{+}\left[\sigma^{-1}(\mathcal{R}_{2,T+1})-A\mathcal{R}_{1,T}\right]\mathcal{R}_{1,T}^{+}$ (4)
assuming that $\sigma$ is invertible and $\text{rank}(B)=n_{\text{in}}$.
@@
Using (3) and Corollary 1. ∎
The additional regularity condition for matrix $\mathcal{R}_{1,T}$ yields an alternative form of Theorem 1, as follows:
@@
The least squares solution for IR in ESNs is expressed as:
$W_{\mathcal{R}}:=B^{+}\left[\sigma^{-1}(\mathcal{R}_{2,T+1})\mathcal{R}_{1,T}^{+}-A\right]$ (5)
assuming that $\sigma$ is invertible, $\text{rank}(B)=n_{\text{in}}$, and $\text{rank}(\mathcal{R}_{1,T})=n_{\text{r}}$.
@@
Using Theorem 1 and $\mathcal{R}_{1,T}\mathcal{R}_{1,T}^{+}=I$. ∎
Theorems 1 and 2 show that the readout layer can be trained to solve IR without relying on the original input $d_{t}$; instead, the fixed ESN properties $\sigma$, $A$, and $B$ can be used. Hence, IR in ESNs can be formalized as UL rather than SL.
@@ 2.3.2 Remarks
Using Theorem 2, the least squares solution of IR in ESNs is expressed as $W_{\mathcal{R}}=B^{+}(\hat{A}_{T}-A)$, where $\hat{A}_{T}$ minimizes the following square loss function:
$L_{\sigma,T}(\hat{A}):=\sum_{t=1}^{T-1}\|\sigma^{-1}(r_{t+1})-\hat{A}r_{t}\|^{2}.$ (6)
Note that $\hat{A}_{T}$ can be calculated without using original inputs $d_{t}$, $A$, and $B$. Apart from using the Moore–Penrose inverse of a matrix, numerous optimizations are available to obtain the minimizer of the loss function $L_{\sigma,T}(\hat{A})$. For example, the recursive least square (RLS) algorithm (Haykin, 2014) sequentially updates $(\hat{A}_{t},P_{t})$ to $(\hat{A}_{t+1},P_{t+1})$ based on newly obtained data $(r_{t},\sigma^{-1}(r_{t+1}))$, where $P_{T}$ is the recursive estimation of the precision matrix of the reservoir state. Hence, a sequential algorithm for IR in ESNs based on UL is presented (Algorithm 1).

### v4

@@ Abstract
Echo state networks (ESNs) are a class of recurrent neural networks in which only the readout layer is trainable, while the recurrent and input layers are fixed. This architectural constraint enables computationally efficient processing of time-series data. Traditionally, the readout layer in ESNs is trained using supervised learning with target outputs. In this study, we focus on input reconstruction (IR), where the readout layer is trained to reconstruct the input time series fed into the ESN. We show that IR can be achieved through unsupervised learning (UL), without access to supervised targets, provided that the ESN parameters are known a priori and satisfy invertibility conditions. This formulation allows applications relying on IR, such as dynamical system replication and noise filtering, to be reformulated within the UL framework via straightforward integration with existing algorithms. Our results suggest that prior knowledge of ESN parameters can reduce reliance on supervision, thereby establishing a new principle: not only by fixing part of the network parameters but also by exploiting their specific values. Furthermore, our UL-based algorithms for input reconstruction and related tasks are suitable for autonomous processing, offering insights into how analogous computational mechanisms might operate in the brain in principle. These findings contribute to a deeper understanding of the mathematical foundations of ESNs and their relevance to models in computational neuroscience.
@@ 1 Introduction
Reservoir computing (RC) is an efficient computational framework for processing time-series data. Models such as the echo state network (ESN) (Jaeger,, 2001) and the liquid state machine (Maass et al.,, 2002) share a common architecture: fixed input and reservoir layers, and a trainable readout layer (see, e.g., Tanaka et al.,, 2019; Nakajima,, 2020, for reviews); see also Figure 1(a). In standard RC settings, the readout layer is trained to reproduce a target time series, which must be defined based on the task (see, e.g., Schrauwen et al.,, 2007; Lukoševičius et al.,, 2012; Zhang and Vargas,, 2023; Yan et al.,, 2024). In certain cases (e.g., Lu et al.,, 2018; Vlachas et al.,, 2020; Bollt,, 2021), the input time series itself serves as the target. In this paper, we refer to such tasks as input reconstruction.
@@
Input reconstruction (IR) is one of the simplest yet most fundamental tasks in RC. The RC performance is based on two time-series processing characteristics: preservability and transformability of the input time series within RC (Maass et al.,, 2002; Jaeger,, 2002; Dambre et al.,, 2012). IR, by definition, requires preserving the input without transformation; consequently, it is the most basic time-series processing task in RC. Despite its simplicity, IR in RC is sufficient for learning a dynamical system that generates an input time series (Lu et al.,, 2017, 2018; Röhm et al.,, 2021; Grigoryeva et al.,, 2023). Here, we focus on IR to investigate the fundamental properties of RC.
@@
IR has been studied not only within the context of RC but also through two major unsupervised learning approaches: noise filtering (Candy,, 2005) and blind source separation (Ghahramani,, 2003). Noise filtering methods, such as the Kalman filter (Kalman,, 1960) and its nonlinear extensions (see Daum,, 2005, for a review), rely on prior knowledge about a system model governing the input time series. In noise filtering, this prior knowledge is used to reconstruct the original input time series or system state using noisy observations (Figure 1(b)). In contrast, blind source separation requires no explicit system model and instead exploits statistical properties of the signals. Examples include principal component analysis (Pearson,, 1901), independent component analysis (Jutten and Herault,, 1991), and slow feature analysis (Wiskott and Sejnowski,, 2002). Compared to noise filtering, blind source separation typically provides approximate IR by recovering the permuted and scaled versions of the original inputs. This method relies on minimal assumptions, such as invertibility of mixing transformations, and aims to infer unobserved source signals from their observed mixtures (Figure 1(c)).
@@
Despite the conceptual relevance of IR to unsupervised learning (UL), the standard approach in RC remains supervised learning (SL) of the readout layer (Figure 1(a)). Several studies have explored combining RC with noise filtering (e.g., Han and Wang,, 2009; Tsai et al.,, 2010; Tomizawa and Sawada,, 2021; Goswami et al.,, 2021) or blind source separation (e.g., Antonelo and Schrauwen,, 2009; Dinh and Hinaut,, 2020; Steiner et al.,, 2023). However, in most of these approaches, the unobserved original input time series is still used as supervised data for training the readout layer. Although the above list is not exhaustive, this reliance on supervised signals remains widespread. From the perspective of unsupervised IR, the prevailing problem formulation in RC seems conceptually inconsistent with the foundational goals of noise filtering and blind source separation, which do not assume supervision for IR.
@@
The objective of this study is to refine and advance the problem formulation of IR in RC. We propose a novel formulation of IR based on UL. Specifically, we aim to replace the conventional assumption of access to the unobserved original input time series with prior knowledge of RC parameters that satisfy certain conditions (Figure 1(d)). This formulation enables IR to be defined independently of any input time series, thereby serving as a standard formulation of IR within the RC framework. To illustrate the utility of this perspective, we develop algorithms for dynamical system replication and noise filtering that do not explicitly rely on the original input time series. In addition, we conduct numerical experiments to demonstrate the feasibility and validity of the proposed approach.
@@
This paper is organized as follows. In Section 2, we introduce ESN as the theoretical model of RC, define the IR task, and propose its reformulation within an UL framework. In Section 3, we demonstrate that dynamical system replication and noise filtering in ESNs can be theoretically achieved via IR, thus reformulated as UL tasks. Additionally, we provide numerical examples illustrating both tasks solved using the proposed method. In Section 4, we discuss the implications of our findings, highlighting how IR, dynamical system replication, and noise filtering in ESNs are fundamentally redefined under the UL framework. Finally, we present our conclusions in Section 5.
@@ 2 Unsupervised Input Reconstruction
In this section, we formalize UL for IR in RC. Initially, we introduce ESN (Jaeger,, 2001) as a numerically implementable and mathematically tractable RC model. Next, we formalize IR in ESNs. Finally, we characterize the mathematical conditions under which UL becomes feasible for IR in ESNs.
@@ 2.1 Echo State Network
$r_{t+1}=g(d_{t},r_{t}):=\sigma\left(Ad_{t}+Br_{t}\right),$ (2.1)
where $\sigma:\mathbb{R}^{n_{\text{r}}}\to\mathbb{R}^{n_{\text{r}}}$, $A\in\mathbb{R}^{n_{\text{r}}\times n_{\text{in}}}$, and $B\in\mathbb{R}^{n_{\text{r}}\times n_{\text{r}}}$. The activation function $\sigma$ and the weight matrices $A$ and $B$ are fixed for each ESN; they are referred to as the activation, input, and reservoir layers, respectively.
@@
The standard approach to learning $W$ is the least squares method, which requires the desired output data $o_{t}$. The least squares solution for the readout layer parameter that minimizes $\|Wr_{t}-o_{t}\|^{2}$ is
$W_{\mathcal{O}}:=\mathcal{O}_{1,T}\mathcal{R}_{1,T}^{+}\in\mathbb{R}^{n_{\text{out}}\times n_{\text{r}}},$ (2.2)
@@ 2.2 Input Reconstruction
In IR, we consider the case where the desired ESN output is equal to its input, i.e., $o_{t}=d_{t}\;(n_{\text{out}}=n_{\text{in}})$ for all time steps $t$. Then, using the standard approach introduced in Equation 2.2, the solution for IR is
$W_{\mathcal{D}}:=\mathcal{D}_{1,T}\mathcal{R}_{1,T}^{+}\in\mathbb{R}^{n_{\text{out}}\times n_{\text{r}}}.$ (2.3)
Notation $W_{\mathcal{D}}$ emphasizes that its calculation requires samples obtained from the original input time series $\mathcal{D}_{1,T}$. Thus, the standard approach for IR in ESN assumes that samples obtained from the original input time series $\mathcal{D}_{1,T}$ are available during the training of the readout layer. Nevertheless, the objective remains to reconstruct the original input during testing. We address the distinction in the problem settings between training and testing by introducing a training method for the readout layer that does not explicitly rely on the original inputs $\mathcal{D}_{1,T}$.
@@ 2.3 Unsupervised Learning for Input Reconstruction
The most vital key point of our method is that the original input $d_{t}$ can be expressed using reservoir state $r_{t}$ under certain invertibility assumptions. Specifically, we have the following lemma:
@@
$d_{t}=A^{+}\left[\sigma^{-1}(r_{t+1})-Br_{t}\right]$ for all $t\in\mathbb{Z}$ if $\sigma$ is invertible and $\text{rank}(A)=n_{\text{in}}$.
@@
Solving the ESN state update Equation 2.1 $r_{t+1}=\sigma\left(Ad_{t}+Br_{t}\right)$ for $d_{t}$ yields:
$r_{t+1}$ $=\sigma\left(Ad_{t}+Br_{t}\right)$
$\sigma^{-1}(r_{t+1})$ $=Ad_{t}+Br_{t}$ $(\because\sigma\text{ is invertible})$
$\sigma^{-1}(r_{t+1})-Br_{t}$ $=Ad_{t}$
$A^{+}[\sigma^{-1}(r_{t+1})-Br_{t}]$ $=d_{t}$ $(\because A\text{ has full column rank})$
@@
$\mathcal{D}_{1,T}=A^{+}\left[\sigma^{-1}(\mathcal{R}_{2,T+1})-B\mathcal{R}_{1,T}\right]$ for all $t\in\mathbb{Z}$ if $\sigma$ is invertible and $\text{rank}(A)=n_{\text{in}}$.
@@
Assume that the activation function $\sigma$ is invertible, and $\operatorname{rank}(A)=n_{\text{in}}$. Then, the least-squares solution for IR in ESNs is given by
$W_{\mathcal{R}}=A^{+}\left[\sigma^{-1}(\mathcal{R}_{2,T+1})-B\mathcal{R}_{1,T}\right]\mathcal{R}_{1,T}^{+}.$ (2.4)
@@
From Corollary 1, we have the identity $\mathcal{D}_{1,T}=A^{+}\left[\sigma^{-1}(\mathcal{R}_{2,T+1})-B\mathcal{R}_{1,T}\right],$ which expresses the input time series explicitly in terms of the ESN parameters and internal states. Substituting this expression into the least-squares solution for input reconstruction 2.3, $W_{\mathcal{D}}=\mathcal{D}_{1,T}\mathcal{R}_{1,T}^{+},$ yields the desired formulation of the readout matrix $W_{\mathcal{R}}$ for unsupervised input reconstruction. ∎
The additional regularity condition for matrix $\mathcal{R}_{1,T}$ yields an alternative form of Theorem 1, stated below.
@@
Assume that the activation function $\sigma$ is invertible, $\operatorname{rank}(A)=n_{\text{in}}$, and $\operatorname{rank}(\mathcal{R}_{1,T})=n_{\text{r}}$. Then, the least-squares solution for IR in ESNs is given by
$W_{\mathcal{R}}=A^{+}\left[\sigma^{-1}(\mathcal{R}_{2,T+1})\mathcal{R}_{1,T}^{+}-B\right].$ (2.5)
@@
Since $\operatorname{rank}(\mathcal{R}_{1,T})=n_{\text{r}}$, we have the identity $\mathcal{R}_{1,T}\mathcal{R}_{1,T}^{+}=I$. Substituting this into the expression for $W_{\mathcal{R}}$ in Theorem 1 yields Equation 2.5. ∎
Theorems 1 and 2 show that the readout layer can be trained to solve IR without explicitly relying on the original input $d_{t}$; instead, the fixed ESN properties $\sigma$, $A$, and $B$ can be used. Hence, IR in ESNs can be formalized as UL, not SL.
Our theory of UL in ESN for IR relies on three independent regularity conditions: (i) invertibility of the activation function $\sigma$, (ii) full column rank of the input weight matrix $A$, and (iii) full row rank of the reservoir state matrix $\mathcal{R}_{1,T}$.
Conditions (i) and (ii) are required to transform the readout matrix from $W_{\mathcal{D}}$ (Equation 2.3) to $W_{\mathcal{R}}$ (Equation 2.4). This transformation involves reconstructing the ESN input sequence $\mathcal{D}_{1,T}$, which is used as targets in supervised training, from the reservoir states $\mathcal{R}_{1,T+1}$ (see Corollary 1). Violations of these conditions result in incomplete input recovery and degraded IR performance (Figure S2).
In practice, condition (ii) is generally satisfied, since random matrices drawn from distributions with positive variance are almost surely full rank. In contrast, condition (i) is more critical: commonly used activation functions such as ReLU are not globally invertible. Nevertheless, we numerically confirmed that IR remains feasible under ReLU, albeit with reduced precision. Despite the loss in accuracy, the reconstruction still performs well above chance (Figure S1, S2), indicating meaningful reconstruction capability in non-invertible settings. See Supplementary Material for details.
Condition (iii) is necessary to derive the simplified form of the readout matrix (Equation 2.5) from the general expression (Equation 2.4). This condition may fail when the input lacks complexity or when the reservoir states exhibit insufficient diversity due to the echo state property. However, we found that injecting noise into the ESN input improves the effective rank of $\mathcal{R}_{1,T}$ and mitigates this issue (Figure S3). When conditions (i) and (ii) are satisfied but condition (iii) is not, the SL solution $W_{\mathcal{D}}$ and the intermediate UL solution $W_{\mathcal{R}}$ (Equation 2.4) remain mathematically equivalent, while $W_{\mathcal{R}}$ (Equation 2.5) deviates. Importantly, this deviation does not necessarily imply performance degradation. We observed that in such ill-conditioned cases, the UL solution $W_{\mathcal{R}}$ (Equation 2.5) outperforms its intermediate counterpart (Equation 2.4), which is guaranteed to match the SL solution (Figure S3). This highlights the numerical advantage of the proposed UL-based IR framework. Additional results are provided in the Supplementary Material.
@@ 2.3.1 Remarks
Theorem 2 reveals that the least squares solution of IR in ESNs can be written as $W_{\mathcal{R}}=A^{+}(\hat{B}_{T}-B)$, where $\hat{B}_{T}$ minimizes the following square loss function:
$L_{\sigma,T}(\hat{B}):=\sum_{t=1}^{T-1}\|\sigma^{-1}(r_{t+1})-\hat{B}r_{t}\|^{2}.$ (2.6)
Note that $\hat{B}_{T}$ can be calculated without using original inputs $d_{t}$, $A$, and $B$. Apart from using the Moore–Penrose inverse of a matrix, numerous optimizations are available to obtain the minimizer of the loss function $L_{\sigma,T}(\hat{B})$. For example, the recursive least square (RLS) algorithm (Haykin,, 2002) sequentially updates $(\hat{B}_{t},P_{t})$ to $(\hat{B}_{t+1},P_{t+1})$ based on newly obtained data $(r_{t},\sigma^{-1}(r_{t+1}))$, where $P_{T}$ is the recursive estimation of the precision matrix of the reservoir state. Hence, a sequential algorithm for IR in ESNs based on UL is presented (Algorithm 1).